OpenStudy (wcrmelissa2001):

Help please thanks :DIf the y-axis is tangential to the curve px^2+y^2-2px+2qy+r=0, fin r in terms of q. If this term does not meet the x-axis, find the range of values of p in terms of q

OpenStudy (welshfella):

At the point where y is tangential to the curve x = 0 right?

OpenStudy (wcrmelissa2001):

yup that part i got

OpenStudy (wcrmelissa2001):

i also tried equating discriminant and this but hm (thanks again)

OpenStudy (welshfella):

- by the way - the last question about the parabola the value of a is -0.5

OpenStudy (wcrmelissa2001):

ACK WAIT WHAT i got 5..lemme check

OpenStudy (welshfella):

no need to use the vertex form

OpenStudy (wcrmelissa2001):

i keyed in my answer and it's right. like i check with online grapher

OpenStudy (welshfella):

what a = 5?

OpenStudy (wcrmelissa2001):

at least the intercepts are

OpenStudy (welshfella):

a must be negative as the graph has a maximum

OpenStudy (welshfella):

check the post

OpenStudy (wcrmelissa2001):

oh no. what is this. :( is it not 5. damn. yeah it has to be negative what am i doing. sorry i'lll go check. meanwhile you mind helping me with this qn?

OpenStudy (wcrmelissa2001):

omg i get it (the other qn) but i don't get why i'm wrong using the vertex method

OpenStudy (welshfella):

I am trying I had an operation yesterday and i'm not at my best.

OpenStudy (wcrmelissa2001):

ack! Rest well. Don't need to help me haha just go and rest :D

OpenStudy (welshfella):

Yes i think i will

OpenStudy (wcrmelissa2001):

:D

OpenStudy (wcrmelissa2001):

@phi (if you want to solve this one that is)

OpenStudy (baru):

ok we have \[px^2 + y^2 -2px +2qy +r=0\] we know that the y axis is tangent which means y' at x=0 is infinity differentiate \[2px +2yy' -2p +2qy' =0 \\ px +y'(y+q) -2p =0\\at~x=0\\y' =\frac{2p}{y+q}\]

OpenStudy (baru):

we want y' equal to infinity so y=-q so (0,-q) is a point on the graph, substitute \[ q^2 -2q^2 +r=0\\q^2=r\]

OpenStudy (wcrmelissa2001):

wait what is y'

OpenStudy (baru):

y' is dy/dx i just wrote y' because its a neater notation

OpenStudy (baru):

making sense now?

OpenStudy (baru):

for the second part, we know when y=0, x is not zero so sub y=0 in the equation to get \[px^2 -2px +r=0\] we know that r=q^2 \[px^2 -2px +q^2=0\]

OpenStudy (baru):

i dont know where to go from here, perhaps complete the square for x^2 terms?

OpenStudy (wcrmelissa2001):

sorry what's dy/dx?

OpenStudy (baru):

the derivative of y with respect to x

OpenStudy (welshfella):

y' gives the slope of the tangent to the line. The y-axis was the the tangent - a vertical line whose slope is infinite.

OpenStudy (baru):

@welshfella does the \(q^2=r\) calculations look correct? any idea how to proceed for the second part of the question?

OpenStudy (welshfella):

yes . Looks correct. No. I'm not sure how to proceed with part 2.

OpenStudy (welshfella):

'If this term does not meet the x axis' is confusing Shouldn't it be 'If the curve'

OpenStudy (baru):

yea, i just assumed thats what it meant

OpenStudy (baru):

ok we have \[px^2 -2px +q^2=0\] this equation should not have x=1 as root, so we do the whole b^2-4ac thing ?

OpenStudy (baru):

*x=0 as a root

OpenStudy (baru):

i just get q not equal to zero...

OpenStudy (welshfella):

why not use completing the square?

OpenStudy (baru):

yea \[(x-1)^2 = \frac{p-q^2}{p}\] so \[(0-1)^2 \neq \frac{p-q^2}{p}\]

OpenStudy (baru):

so again, just q \(\neq\)0 ?

OpenStudy (welshfella):

yea I see what you mean . I cant see any problem with that.

OpenStudy (welshfella):

To be honest I always struggle a bit with this type of problem.

OpenStudy (baru):

oh well.. i guess we'll leave it at this. @wcrmellissa2001 will have to verify. Hope you recover quickly @welshfella , good luck!

OpenStudy (welshfella):

yes thank you . I'm still not at full strength.

OpenStudy (baru):

and yet you cant keep away from openstudy can you :P it's an addiction

OpenStudy (welshfella):

right!!

OpenStudy (welshfella):

This curve is an ellipse how about this (x - 1)^2 = p - q^2 ------ p when x = 1 p - q^2 ------- = 0 p and the curve touches the x-axis at this point p = q^2 at x = 1 any value of q^2 greater than p would make (x - 2)^2 negative so for p > x^2 the curve will be lower than the x-axis rhat's the range they want

OpenStudy (welshfella):

https://www.desmos.com/calculator/7varuvy9np

OpenStudy (wcrmelissa2001):

THANKS GUYS :D

OpenStudy (wcrmelissa2001):

maybe my teacher can clarify

OpenStudy (welshfella):

yes i'm pretty sure the above is right though q^2 cant be greater than p because a square cant be negative and as you see from the graph when p = q^2 the curve just touches the axis when p > q^2 the whole curve is below the x axis.

OpenStudy (wcrmelissa2001):

ahh i actually sort of get it. I'm so proud. REST WELL @WELSHFELLA!!

OpenStudy (welshfella):

yes OK I'm sure you've learned something today.

OpenStudy (baru):

yes!! what was i thinking lol if it never intersects the x axis then this \[(x-1)^2 = \frac{p-q^2}{p}\] should have no real solutions nice one @welshfella

OpenStudy (wcrmelissa2001):

i thnik it's something to do with discriminants for that one

OpenStudy (welshfella):

yea for there to be 2 real roots D > 0 one root D = 0 and no roots (doesn't pas the x-axis ) D < 0.

OpenStudy (welshfella):

lets try that

OpenStudy (welshfella):

D = 4p(^2 - 4pq^2 < 0 for the graph not to touch the x axis 4p(p - q^2) < 0 this gives p < q^2 - the opposite to what i said earlier!!! I'll draw some more graphs to confirm this

OpenStudy (welshfella):

https://www.desmos.com/calculator/nr07exibs6 check this one out p = 8 and q^2 = 4 so p> q^2 is wrong I'll try one with p < q^2

OpenStudy (welshfella):

ttps://www.desmos.com/calculator/wjlgertejj Yes this is not touching the x axis that confirms p < q^2

OpenStudy (welshfella):

There must have been a flaw in my argument which showed p > q^2 but I cant see it. You were right @wcrmelissa2001 - the discriminant was the way to go!

OpenStudy (wcrmelissa2001):

Yup haha just so happened I checked with my classmate before I saw this... method is same as yours! Thanks :D :D